Sheridan Company purchased a new machine on October 1, 2022, at a cost of $104,100. The company estimated that the machine has a salvage value of $7,140. The machine is expected to be used for 96,000 working hours during its 10-year life.

Compute depreciation using the following methods in the year indicated.

Declining-balance using double the straight-line rate for 2022 and 2023.

2022

2023

Depreciation using the Declining-balance method

$enter a dollar amount

$enter a dollar amount

eTextbook and Media

Calculate the depreciation cost per hour. (Round answer to 2 decimal places, e.g. $1.25)

Depreciation cost per hour

$enter the depreciation cost per hour in dollars rounded to 2 decimal places


Units-of-activity for 2022, assuming machine usage was 500 hours. (Round answer to 0 decimal places, e.g. 5,275.)

Depreciation using the Units-of-activity method for 2022

$enter the units-of-activity depreciation cost for 2022 in dollars
Answer #1

Double declining rate = 200/useful life = 200/10 = 20%

2022 2023
Double declining

= 104,100*20%*3/12

= 5205

= (104,100-5205)*20%

= 19,779

.

Depreciation cost per hour

(104,100-7,140)/96,000

1.01

.

Depreciation

(1.01*500)

505